# BioT5+: Towards Generalized Biological Understanding with IUPAC Integration and Multi-task Tuning

### v1

@@ Abstract
Recent research trends in computational biology have increasingly focused on integrating text and bio-entity modeling, especially in the context of molecules and proteins. However, previous efforts like BioT5 faced challenges in generalizing across diverse tasks and lacked a nuanced understanding of molecular structures, particularly in their textual representations (e.g., IUPAC). This paper introduces BioT5+, an extension of the BioT5 framework, tailored to enhance biological research and drug discovery. BioT5+ incorporates several novel features: integration of IUPAC names for molecular understanding, inclusion of extensive bio-text and molecule data from sources like bioRxiv and PubChem, the multi-task instruction tuning for generality across tasks, and a novel numerical tokenization technique for improved processing of numerical data. These enhancements allow BioT5+ to bridge the gap between molecular representations and their textual descriptions, providing a more holistic understanding of biological entities, and largely improving the grounded reasoning of bio-text and bio-sequences. The model is pre-trained and fine-tuned with a large number of experiments, including 3 types of problems (classification, regression, generation), 15 kinds of tasks, and 21 total benchmark datasets, demonstrating the remarkable performance and state-of-the-art results in most cases. BioT5+ stands out for its ability to capture intricate relationships in biological data, thereby contributing significantly to bioinformatics and computational biology. Our code is available at https://github.com/QizhiPei/BioT5.
@@
Qizhi Pei1, Lijun ${}^{\ast}$, Kaiyuan Gao3, Xiaozhuan Liang4, Yin Fang4, Jinhua Zhu5, Shufang Xie1, Tao Qin2, Rui Yan1thanks:   Corresponding authors: Lijun Wu (lijuwu@microsoft.com) and Rui Yan (ruiyan@ruc.edu.cn) 1Gaoling School of Artificial Intelligence, Renmin University of China 2Microsoft Research  3Huazhong University of Science and Technology 4Zhejiang University  5University of Science and Technology of China {qizhipei,shufangxie,ruiyan}@ruc.edu.cn {lijuwu,taoqin}@microsoft.comim_kai@hust.edu.cn {liangxiaozhuan,fangyin}@zju.edu.cnteslazhu@mail.ustc.edu.cn
@@ 1 Introduction
Molecules and proteins are two crucial bio-entities in drug discovery, forming the foundation of biological activities (Dara et al., 2022). A molecule can be represented by its SMILES (Weininger, 1988; Weininger et al., 1989) or SELFIES (Krenn et al., 2020) sequence, and a protein can be described by a FASTA (Lipman and Pearson, 1985; Pearson and Lipman, 1988) sequence. With the advancement of Language Models (LMs), an increasing body of work focuses on understanding the molecules and proteins by modeling their bio-sequences (Chithrananda et al., 2020; Rives et al., 2021; Lin et al., 2022).
Notably, biological literature (Canese and Weis, 2013; White, 2020) is full of extensive information on molecules and proteins. When a biological entity is mentioned in such literature, its context is predominantly centered around a description of some characteristics of the entity. Consequently, there has been a growing body of work dedicated to the joint modeling of text and biological entities, such as Galactica (Taylor et al., 2022), MolXPT (Liu et al., 2023c), BioT5 (Pei et al., 2023) and BioMedGPT (Luo et al., 2023c), which are all scientific models trained on text, molecule and protein sequences. Despite their achievements, substantial opportunities for enhancement still remain: (i) Prior works neglect the importance of modeling the textual name of molecules, such as International Union of Pure and Applied Chemistry (IUPAC), which provides a standard and systematic naming method for ensuring uniformity and clarity across the scientific community. Different from SMILES and SELFIES, IUPAC bears a closer resemblance to natural language that is evident in its widespread adoption within scientific literature (Klinger et al., 2008). (ii) Previous models were predominantly specialist models, necessitating the training of a separate model for each downstream task, thereby lacking in generality and increasing the training and developing cost Liu et al. (2023c); Pei et al. (2023). (iii) Most of the previous models based on T5 (Raffel et al., 2020) and GPT (Brown et al., 2020) architectures only focus on the classification tasks since they do not implement specialized tokenization for numerical data, which results in their suboptimal adaptation to regression tasks.
@@
(3) Multitask Instruction Tuning: BioT5+ employs multi-task instruction tuning strategy for downstream tasks rather than the separate specialized model training for each task. By leveraging a unified and multi-task training framework, BioT5+ can seamlessly integrate knowledge from diverse tasks, enhancing its predictive power and generalization capabilities across different biological and chemical domains.
@@ 2.1 Biological Cross-modal Models
Recent advancements in LLMs have led to an increased focus on jointly modeling molecules, proteins, and text, aiming to enhance the understanding of bio-entities through textual data.
@@ 3.1 Pre-training Corpus
The pre-training corpus consists of 4 classes: (1) Single-modal data, including molecule SELFIES with IUPAC name from PubChem (Kim et al., 2019), molecule SELFIES from ZINC20 (Irwin et al., 2020), protein FASTA from Uniref50 (Suzek et al., 2007), and general text from “Colossal Clean Crawled Corpus” (C4) (Raffel et al., 2020). For the molecule from PubChem, we concatenate the IUPAC name and SELFIES for pre-training as shown in Figure 3. (2) Wrapped text, where the molecule or gene/protein names are suffixed with corresponding FASTA sequence representation. We employ BERN2 (Sung et al., 2022), a neural-based Named Entity Recognition (NER) system in the biological domain, to detect and classify occurrences of molecules and proteins within the abstracts of PubMed (White, 2020) and bioRxiv (Sever et al., 2019). For the molecule name, we first standardize the name to its IUPAC name and then append the corresponding SELFIES. For the gene/protein name, we will directly append its FASTA sequence. For the generation of high-quality wrapped text, we also analyze the confidence score distribution predicted by BERN2 (Sung et al., 2022). Only those entities with higher confidence scores were retained to ensure the accuracy and relevance of the appended sequence data. Further detailed descriptions of this process are provided in Appendix Section C. (3) Bio-text, including PubMed (White, 2020) Central full text articles, and bio-texts from PubMed (White, 2020) abstracts and bioRxiv (Sever et al., 2019) abstracts that do not yield identifiable named entities in (2). (4) Molecule-description pairs and protein-description pairs. The molecule-text data is collected from PubChem (Kim et al., 2019) and we also add the IUPAC name to the text description. All molecules and proteins that exist in the downstream Mol-Instructions dataset (Fang et al., 2023) and ChEBI-20 (Edwards et al., 2022) are excluded to prevent data leakage. The protein-text data is the same as BioT5 (Pei et al., 2023).
Remarkable Difference. The primary distinctions between BioT5+ and BioT5 are as follows: (i) BioT5+ integrates IUPAC in the molecular pre-training data, encompassing IUPAC names combined with SELFIES, wrapped text, and molecule-text translation data. More details are in Appendix Section D. (ii) BioT5+ incorporates a broader spectrum of high-quality data, including IUPAC names and SELFIES from PubChem, as well as comprehensive articles from bioRxiv and PubMed Central.
@@ 3.3 Model and Training
Model architecture. BioT5+ follows the T5-v1.1-base [^1] https://huggingface.co/docs/transformers/model_doc/t5v1.1 configuration with vocab size $35,076$ and $252$M parameters.
@@
Multi-task Instruction-based Fine-tuning. After the comprehensive pre-training phase, BioT5+ undergoes multi-task instruction-based fine-tuning. Unlike BioT5 each downstream task has a specialized fine-tuned model, we follow Fang et al., 2023 and  Cao et al., 2023 to categorize downstream tasks and conduct multi-task instruction tuning, which not only saves the repeated tuning cost but also eases the model deployment for the evaluation of multiple tasks. The relevant groupings and information about benchmark tasks and datasets are illustrated in Figure 1, which is simply split by the domains, e.g., molecule-oriented or protein-oriented tasks. This methodology serves a dual purpose: firstly, it bridges the gap between the pre-training and fine-tuning phases, ensuring a smoother transition and integration of learned capabilities. Secondly, it activates and harnesses the general capabilities of BioT5+ across various tasks, demonstrating its versatility and adaptability in handling diverse biological problems.
@@ 4 Experiments and Results
As shown in Figure 1, BioT5+ is extensively evaluated across 21 well-established downstream benchmark datasets, which can be classified into 7 molecule-oriented tasks and 8 protein-oriented tasks with 3 types of problems: classification, regression, and generation. Following Fang et al., 2023, we categorize downstream tasks into different groups for multi-task instruction tuning in the same way, and details about the downstream datasets and baselines are in Appendix Section E.
@@ 4.1 Molecule-oriented Tasks
4.1.1 Molecule Property Prediction
Table 1: Performance (AUROC) comparison on molecule property prediction tasks (classification) on MoleculeNet Wu et al. (2018) benchmark (Best, Second Best). $*{}^{*}$ means LoRA Hu et al. (2022) tuning.
@@ 4.1.1 Molecule Property Prediction
$*{}^{*}$ 68.3 60.1 58.1 -
$*{}^{*}$ 74.8 65.6 62.3 -
InstructMol-G 85.9 64.0 74.0 -
InstructMol-GS 82.3 70.0 68.9 -
@@
Table 2: Performance comparison on chemical reaction-related tasks (Best, Second Best). *** means LoRA tuning.
Model Exactnormal-$\uparrow$ BLEUnormal-$\uparrow$ Levenshteinnormal-$\downarrow$ RDK FTSnormal-$\uparrow$ MACCS FTSnormal-$\uparrow$ Morgan FTSnormal-$\uparrow$ Validitynormal-$\uparrow$
@@
Llama 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Galactica 0.000 0.141 30.760 0.036 0.127 0.051 0.995
Text+Chem T5 0.000 0.225 49.323 0.039 0.186 0.052 0.313
Mol-Instructions 0.044 0.224 23.167 0.237 0.364 0.213 1.000
$*{}^{*}$(LoRA) 0.000 0.283 53.510 0.136 0.294 0.106 1.000
InstructMol-G 0.070 0.890 24.732 0.469 0.691 0.426 1.000
InstructMol-GS 0.129 0.610 19.664 0.444 0.539 0.400 1.000
@@
Llama 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Galactica 0.000 0.468 35.021 0.156 0.257 0.097 0.946
Text+Chem T5 0.239 0.782 20.413 0.705 0.789 0.652 0.762
Mol-Instructions 0.045 0.654 27.262 0.313 0.509 0.262 1.000
$*{}^{*}$(LoRA) 0.012 0.804 29.947 0.499 0.649 0.407 1.000
InstructMol-G 0.153 0.906 20.155 0.519 0.717 0.457 1.000
InstructMol-GS 0.536 0.967 10.851 0.776 0.878 0.741 1.000
@@
Llama 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Galactica 0.000 0.452 34.940 0.167 0.274 0.134 0.986
Text+Chem T5 0.141 0.765 24.043 0.685 0.765 0.585 0.698
Mol-Instructions 0.009 0.705 31.227 0.283 0.487 0.230 1.000
$*{}^{*}$(LoRA) 0.000 0.283 53.510 0.136 0.294 0.106 1.000
InstructMol-G 0.114 0.586 21.271 0.422 0.523 0.285 1.000
InstructMol-GS 0.407 0.941 13.967 0.753 0.852 0.714 1.000
@@
Table 3: Performance (MAE) comparison on molecule property prediction tasks (regression) on QM9 dataset from MoleculeNet Wu et al. (2018) benchmark (Best, Second Best). $\Delta{\epsilon}$ means HOMO-LUMO gap. $\dagger$ means few-shot in-context learning results.
@@
Mol-Instructions 0.0210 0.0210 0.0203 0.0210
InstructMol-G 0.0060 0.0070 0.0082 0.0070
InstructMol-GS 0.0048 0.0050 0.0061 0.0050
@@
4.1.2 Chemical Reaction-related Tasks
4.1.3 Molecule Description Generation
@@
Model BLEU-2normal-$\uparrow$ BLEU-4normal-$\uparrow$ ROUGE-1normal-$\uparrow$ ROUGE-2normal-$\uparrow$ ROUGE-Lnormal-$\uparrow$ METEORnormal-$\uparrow$
@@
Mol-Instructions 0.249 0.171 0.331 0.203 0.289 0.271
InstructMol-G 0.466 0.365 0.547 0.365 0.479 0.491
InstructMol-GS 0.475 0.371 0.566 0.394 0.502 0.509
@@ 4.1.4 Description-guided Molecule Design
Model BLEUnormal-$\uparrow$ Exactnormal-$\uparrow$ Levenshteinnormal-$\downarrow$ MACCS FTSnormal-$\uparrow$ RDK FTSnormal-$\uparrow$ Morgan FTSnormal-$\uparrow$ FCDnormal-$\downarrow$ Text2Molnormal-$\uparrow$ Validitynormal-$\uparrow$
@@
GPT-3.5-turbo (zero-shot) 0.489 0.019 52.13 0.705 0.462 0.367 2.05 0.479 0.802
@@
4.2 Protein-oriented Tasks
4.2.1 Protein Description Generation
4.2.2 Protein Property Prediction
4.2.3 Protein-related Interaction Prediction

### v2

@@ Abstract
Recent research trends in computational biology have increasingly focused on integrating text and bio-entity modeling, especially in the context of molecules and proteins. However, previous efforts like BioT5 faced challenges in generalizing across diverse tasks and lacked a nuanced understanding of molecular structures, particularly in their textual representations (e.g., IUPAC). This paper introduces BioT5+, an extension of the BioT5 framework, tailored to enhance biological research and drug discovery. BioT5+ incorporates several novel features: integration of IUPAC names for molecular understanding, inclusion of extensive bio-text and molecule data from sources like bioRxiv and PubChem, the multi-task instruction tuning for generality across tasks, and a numerical tokenization technique for improved processing of numerical data. These enhancements allow BioT5+ to bridge the gap between molecular representations and their textual descriptions, providing a more holistic understanding of biological entities, and largely improving the grounded reasoning of bio-text and bio-sequences. The model is pre-trained and fine-tuned with a large number of experiments, including 3 types of problems (classification, regression, generation), 15 kinds of tasks, and 21 total benchmark datasets, demonstrating the remarkable performance and state-of-the-art results in most cases. BioT5+ stands out for its ability to capture intricate relationships in biological data, thereby contributing significantly to bioinformatics and computational biology. Our code is available at https://github.com/QizhiPei/BioT5.
@@
Qizhi Pei1, Lijun Wu2, Kaiyuan Gao3, Xiaozhuan Liang4, Yin Fang4, Jinhua Zhu5, Shufang Xie1, Tao Qin2, Rui Yan1,6thanks:   Corresponding authors: Lijun Wu (apeterswu@gmail.com) and Rui Yan (ruiyan@ruc.edu.cn) 1Gaoling School of Artificial Intelligence, Renmin University of China 2Microsoft Research  3Huazhong University of Science and Technology 4Zhejiang University  5University of Science and Technology of China 6Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education {qizhipei,shufangxie,ruiyan}@ruc.edu.cnapeterswu@gmail.com im_kai@hust.edu.cn{liangxiaozhuan,fangyin}@zju.edu.cn teslazhu@mail.ustc.edu.cntaoqin@microsoft.com
@@ 1 Introduction
Molecules and proteins are two crucial bio-entities in drug discovery, forming the foundation of biological activities (Dara et al., 2022; AI4Science and Quantum, 2023). A molecule can be represented by its SMILES (Weininger, 1988; Weininger et al., 1989) or SELFIES (Krenn et al., 2020) sequence, and a protein can be described by a FASTA (Lipman and Pearson, 1985; Pearson and Lipman, 1988) sequence. With the advancement of Language Models (LMs), an increasing body of work focuses on understanding the molecules and proteins by modeling their bio-sequences (Chithrananda et al., 2020; Rives et al., 2021; Lin et al., 2022).
Notably, biological literature (Canese and Weis, 2013; White, 2020) is full of extensive information on molecules and proteins. When a biological entity is mentioned in such literature, its context is predominantly centered around a description of some characteristics of the entity. Consequently, there has been a growing body of work dedicated to the joint modeling of text and biological entities Pei et al. (2024), such as Galactica (Taylor et al., 2022), MolXPT (Liu et al., 2023c), BioT5 (Pei et al., 2023) and BioMedGPT (Luo et al., 2023c), which are all scientific models trained on text, molecule and protein sequences. Despite their achievements, substantial opportunities for enhancement still remain: (1) Prior works neglect the importance of modeling the textual name of molecules, such as International Union of Pure and Applied Chemistry (IUPAC), which provides a standard and systematic naming method for ensuring uniformity and clarity across the scientific community. Different from SMILES and SELFIES, IUPAC bears a closer resemblance to natural language that is evident in its widespread adoption within scientific literature (Klinger et al., 2008). (2) Previous models were predominantly specialist models, necessitating the training of a separate model for each downstream task, thereby lacking in generality and increasing the training and developing cost Liu et al. (2023c); Pei et al. (2023). (3) Most of the previous models based on T5 (Raffel et al., 2020) and GPT (Brown et al., 2020) architectures only focus on the classification tasks since they do not implement specialized tokenization for numerical data, which results in their suboptimal adaptation to regression tasks.
@@
(3) Multi-task Instruction Tuning: BioT5+ employs multi-task instruction tuning strategy for downstream tasks rather than the separate specialized model training for each task. By leveraging a unified and multi-task training framework, BioT5+ can seamlessly integrate knowledge from diverse tasks, enhancing its predictive power and generalization capabilities across different biological and chemical domains.
@@ 2.1 Biological Cross-modal Models
Recent advancements in LLMs have led to an increased focus on jointly modeling molecules, proteins, and text, aiming to enhance the understanding of bio-entities through text.
@@ 3 BioT5+ Framework
3.1 Intuition for IUPAC Integration
IUPAC naming system provides a standardized set of rules for naming chemical compounds, which allows for the precise description of molecular structures and their components (functional groups, chains, and rings), making it a cornerstone in chemical nomenclature. Typically, an IUPAC name is constructed from the names of individual molecules’ constituent parts, reflecting their structure. This includes prefixes, infixes, and suffixes that indicate various chemical groups and structural features, providing a comprehensive description of the molecule. For instance, the IUPAC name for Aspirin is “2-acetyloxybenzoic acid”. Here, “2-acetyloxy” refers to an acetoxy group attached to the second carbon of a benzene ring, and “benzoic acid” indicates a benzene ring with a carboxylic acid group. The resemblance of IUPAC names to natural language, coupled with their prevalent use in scientific literature, makes them an ideal candidate for model pre-training. By pre-training the model on literature that includes IUPAC names, BioT5+ can establish a nuanced understanding of the relationship between molecules and various textual descriptions of their chemical properties.
@@ 3.2 Pre-training Corpus
The pre-training corpus consists of 4 classes: (1) Single-modal data, including molecule SELFIES with IUPAC name from PubChem (Kim et al., 2019), molecule SELFIES from ZINC20 (Irwin et al., 2020), protein FASTA from Uniref50 (Suzek et al., 2007), and general text from “Colossal Clean Crawled Corpus” (C4) (Raffel et al., 2020). For the molecule from PubChem, we concatenate the IUPAC name and SELFIES for pre-training as shown in Figure 3. (2) Wrapped text, where the molecule or gene/protein names are suffixed with corresponding sequence representation. We employ BERN2 (Sung et al., 2022), a neural-based Named Entity Recognition (NER) system in the biological domain, to detect and classify occurrences of molecules and proteins within the abstracts of PubMed (White, 2020) and bioRxiv (Sever et al., 2019). For the molecule name, we first standardize the name to its IUPAC name and then append the corresponding SELFIES. For the gene/protein name, we will directly append its FASTA sequence. For the generation of high-quality wrapped text, we also analyze the confidence score distribution predicted by BERN2 (Sung et al., 2022). Only those entities with higher confidence scores were retained to ensure the accuracy and relevance of the appended sequence data. Further detailed descriptions of this process are provided in Appendix Section C. (3) Bio-text, including PubMed (White, 2020) Central full text articles, and bio-texts from PubMed (White, 2020) abstracts and bioRxiv (Sever et al., 2019) abstracts that do not yield identifiable named entities in (2). (4) Molecule-description pairs and protein-description pairs. The molecule-text data is collected from PubChem (Kim et al., 2019) and we also add the IUPAC name to the text description. All molecules and proteins that exist in the downstream Mol-Instructions dataset (Fang et al., 2023) and ChEBI-20 (Edwards et al., 2022) are excluded to prevent data leakage. The protein-text data is the same as BioT5 (Pei et al., 2023).
Remarkable Difference. The primary distinctions between BioT5+ and BioT5 are as follows: (1) BioT5+ integrates IUPAC in the molecular pre-training data, encompassing IUPAC names combined with SELFIES, wrapped text, and molecule-text translation data. More details are in Appendix Section D. (2) BioT5+ incorporates a broader spectrum of high-quality data, including IUPAC names and SELFIES from PubChem, as well as comprehensive articles from bioRxiv and PubMed Central.
@@ 3.4 Model and Training
Model architecture. BioT5+ adopts the same architecture as the BioT5 (Pei et al., 2023), which follows the T5-v1.1-base [^1] https://huggingface.co/docs/transformers/model_doc/t5v1.1 configuration with vocab size $35,076$ and $252$M parameters.
@@
Multi-task Instruction-based Fine-tuning. After the comprehensive pre-training phase, BioT5+ undergoes multi-task instruction-based fine-tuning. Unlike BioT5 where each downstream task has a specialized fine-tuned model, we follow Fang et al., 2023 and  Cao et al., 2023 to categorize downstream tasks and conduct multi-task instruction tuning, which not only saves the repeated tuning cost but also eases the model deployment for the evaluation of multiple tasks. The relevant groupings and information about benchmark tasks and datasets are illustrated in Figure 1, which is simply split by the domains, e.g., molecule-oriented or protein-oriented tasks. This methodology serves a dual purpose: firstly, it bridges the gap between the pre-training and fine-tuning phases, ensuring a smoother transition and integration of learned capabilities. Secondly, it activates and harnesses the general capabilities of BioT5+ across various tasks, demonstrating its versatility and adaptability in handling diverse biological problems.
@@ 4 Experiments and Results
As shown in Figure 1, BioT5+ is extensively evaluated across 21 well-established downstream benchmark datasets, which can be classified into 7 molecule-oriented tasks and 8 protein-oriented tasks with 3 types of problems: classification, regression, and generation. Following Fang et al., 2023, we categorize downstream tasks into different groups for multi-task instruction tuning in the same way, and details about the downstream datasets and baselines are in Appendix Section G.
@@ 4.1 Molecule-oriented Tasks
Table 1: Performance (AUROC) comparison on molecule property prediction tasks (classification) on MoleculeNet Wu et al. (2018) benchmark (Best, Second Best). means LoRA Hu et al. (2022) tuning.
@@
Vicuna-v1.3-7B 68.3 60.1 58.1 -
Llama-2-7B-chat 74.8 65.6 62.3 -
InstructMol-G-6.9B 85.9 64.0 74.0 -
InstructMol-GS-6.9B 82.3 70.0 68.9 -
@@
Table 2: Performance comparison on chemical reaction-related tasks (Best, Second Best). * means LoRA tuning.
[…]
@@
Llama-7B 0.000 0.003 28.040 0.037 0.001 0.001 0.001
Galactica-6.7B 0.000 0.141 30.760 0.036 0.127 0.051 0.995
Text+Chem T5-223M 0.000 0.225 49.323 0.039 0.186 0.052 0.313
Mol-Instructions-7B 0.044 0.224 23.167 0.237 0.364 0.213 1.000
Llama-7B(LoRA) 0.000 0.283 53.510 0.136 0.294 0.106 1.000
InstructMol-G-6.9B 0.070 0.890 24.732 0.469 0.691 0.426 1.000
InstructMol-GS-6.9B 0.129 0.610 19.664 0.444 0.539 0.400 1.000
@@
Llama-7B 0.000 0.020 42.002 0.001 0.002 0.001 0.039
Galactica-6.7B 0.000 0.468 35.021 0.156 0.257 0.097 0.946
Text+Chem T5-223M 0.239 0.782 20.413 0.705 0.789 0.652 0.762
Mol-Instructions-7B 0.045 0.654 27.262 0.313 0.509 0.262 1.000
Llama-7B(LoRA) 0.012 0.804 29.947 0.499 0.649 0.407 1.000
InstructMol-G-6.9B 0.153 0.906 20.155 0.519 0.717 0.457 1.000
InstructMol-GS-6.9B 0.536 0.967 10.851 0.776 0.878 0.741 1.000
@@
Llama-7B 0.000 0.036 46.844 0.018 0.029 0.017 0.010
Galactica-6.7B 0.000 0.452 34.940 0.167 0.274 0.134 0.986
Text+Chem T5-223M 0.141 0.765 24.043 0.685 0.765 0.585 0.698
Mol-Instructions-7B 0.009 0.705 31.227 0.283 0.487 0.230 1.000
Llama-7B(LoRA) 0.000 0.283 53.510 0.136 0.294 0.106 1.000
InstructMol-G-6.9B 0.114 0.586 21.271 0.422 0.523 0.285 1.000
InstructMol-GS-6.9B 0.407 0.941 13.967 0.753 0.852 0.714 1.000
@@
Table 3: Performance (MAE) comparison on molecule property prediction tasks (regression) on QM9 dataset from MoleculeNet Wu et al. (2018) benchmark (Best, Second Best). $\Delta{\epsilon}$ means HOMO-LUMO gap.
@@
Mol-Instructions-7B 0.0210 0.0210 0.0203 0.0210
InstructMol-G-6.9B 0.0060 0.0070 0.0082 0.0070
InstructMol-GS-6.9B 0.0048 0.0050 0.0061 0.0050
@@
4.1.1 Molecule Property Prediction
@@
[…]
@@
Mol-Instructions-7B 0.249 0.171 0.331 0.203 0.289 0.271
InstructMol-G-6.9B 0.466 0.365 0.547 0.365 0.479 0.491
InstructMol-GS-6.9B 0.475 0.371 0.566 0.394 0.502 0.509
@@
[…]
@@
Llama2-7B (0-shot) 0.104 0.000 84.18 0.243 0.119 0.089 42.01 0.148 0.631
GPT-3.5-turbo (0-shot) 0.489 0.019 52.13 0.705 0.462 0.367 2.05 0.479 0.802
@@
4.1.2 Chemical Reaction-related Tasks